Which equation has no solution?
A. y = `1/2`x - 2; y = `1/2`x + 1
B. y = 4x - 2; y = -2x + 1
C. y = x + 1; y = x - 2
D. y = 2x + 1; y = 2x - 2

Answer by stanbon(75887)

Which equation has no solution?
A. y = `1/2`x - 2; y = `1/2`x + 1
B. y = 4x - 2; y = -2x + 1
C. y = x + 1; y = x - 2
D. y = 2x + 1; y = 2x - 2
------
Ans:
A because the slopes are the same, so the lines are parallel.
C for the same reason.
D for the same reason
==========
Cheers,
Stan H.